# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Code Smell Detection

### v2

@@ 4.1.2. SLMs
Overall, GraphCodeBERT proved the most effective in identifying CC and DC, while UnixCoder excelled at detecting CM and FE. In comparison, CodeT5 fine-tuned with PEFT methods performed the worst across all types of code smells. CodeBERT performed similarly to GraphCodeBERT, albeit slightly worse. For example, in CC detection, whether using full fine-tuning or PEFT methods, CodeBERT consistently achieved lower F1-scores and MCC values compared to GraphCodeBERT, with differences ranging from 0.11% to 5.93% and 0.21% to 12.10%, respectively. This is reasonable because GraphCodeBERT is built upon CodeBERT and further incorporates code structure information (i.e., data flow) during the pre-training phase (guo2021graphcodebert, ). Across all SLMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods in the code smell detection task, which requires both method- and class-level code understanding. When fine-tuned with PEFT methods, the MCC values of CodeBERT even became negative, reaching -4.05% for LoRA on CC and -2.66% for $(IA)^{3}$ on DC. This suggests that CodeT5 with PEFT methods struggles to perform effectively in code smell detection, highlighting the limitations of its PEFT adaptation. In contrast, the fully fine-tuned CodeT5 demonstrated acceptable performance in code smell detection. The stark performance gap between CodeT5 with full fine-tuning and PEFT methods could stem from CodeT5’s architecture, which may depend more on full fine-tuning for optimal adaptation to the method-level code smell detection task. Compared to full fine-tuning, PEFT methods like LoRA and $(IA)^{3}$ might not provide sufficient adjustments to effectively adapt CodeT5’s encoder-decoder architecture.
@@ 4.1.3. LLMs
Figure 5. Peak GPU memory usage for detecting CC: (a) SLMs with PEFT methods and Full Fine-Tuning; (b) LLMs with PEFT methods
@@
The detection of CC and CM requires method-level code understanding, whereas identifying FE and DC necessitates class-level analysis (see Section 3.3.1). By comparing the detection results of CC and CM with those of FE and DC, we observe that LMs demonstrate comparable effectiveness at both levels of code understanding. Notably, the models even achieve better performance on class-level code smells, as evidenced by the fact that the best-performing model yields significantly lower results on CC compared to the other smell types.
@@ 4.1.6. Best PEFT Method
Generally, for SLMs, both $(IA)^{3}$ and prefix tuning demonstrate strong performance across different models. For LLMs, $(IA)^{3}$ consistently delivers the best results on the detection of all four types of code smells. To be specific, the most effective PEFT method for SLMs on the task of code smell detection varies depending on the model: for instance, CodeBERT performs optimally with $(IA)^{3}$, whereas UnixCoder tends to achieve better results with prefix tuning. However, for LLMs, $(IA)^{3}$ always outperforms other PEFT methods across all studied code smells. Whether for StarCoderBase Models, Deepseek-Coder Models, or CodeLlama-7B, they all achieve superior performance when fine-tuned using $(IA)^{3}$. Additionally, for LLMs, prompt tuning can be employed as an alternative option. Notably, despite achieving promising results on certain models and code smell types, LoRA exhibits instability in performance and can yield markedly poor effectiveness in some cases. An example of this is using LoRA on StarCoderBase-1B for CM detection. A case in point is the use of LoRA on StarCoderBase-1B for detecting CM. While StarCoderBase-1B achieves a minimum MCC of 54.01% with the other three PEFT methods (i.e., prompt tuning, prefix tuning, and $(IA)^{3}$), its performance drops sharply to 7.01% with LoRA, resulting in a disparity of over 40%.
@@
According to Table 6 and Table 7, we can find that all the four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for code smell detection. For small LMs (excluding CodeT5), prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ generally achieve varying degrees of improvement in F1-scores and MCC values compared to full fine-tuning for most code smell types. However, for FE, all PEFT methods perform significantly worse than full fine-tuning. These results suggest that, in most cases, PEFT methods can match or surpass the performance of standard full fine-tuning for code smell detection while adjusting fewer number of trainable parameters.
PEFT methods focus on enhancing the effectiveness of LMs to levels similar to full fine-tuning, while aiming to minimize the use of computational resources, such as peak GPU memory consumption. Taking CC detection as an example, Figure 5 (a) depicts the peak GPU memory consumption during the fine-tuning of the four SLMs. It can be observed that juxtaposed with full fine-tuning, all of the four PEFT methods lead to a substantial reduction in peak GPU memory usage, which further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
Answer to RQ1: Overall, LLMs perform slightly better than SLMs in code smell detection, except for DC. Among different PEFT methods, both $(IA)^{3}$ and prefix tuning are most effective for SLMs, while $(IA)^{3}$ consistently yields the best results for LLMs. All PEFT techniques outperform or match the performance of full fine-tuning for several code smell types, while also reducing peak GPU memory usage, demonstrating their efficiency. However, for the detection of FE, full fine-tuning remains more effective.
@@ 4.2.1. Experiment Setup
Figure 6. MCC results of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA for code smell detection under different rank settings: (a) CC; (b) CM; (c) FE; (d) DC
@@ 4.3. RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
4.3.1. Experiment Setup
To simulate low-resource scenarios, following previous studies (wang2023one, ; liu2023empirical, ), we randomly selected 50, 100, 250, and 500 samples for CC, CM, and DC, respectively, based on the size of our dataset. Note that each subset is balanced, containing an equal number of positive and negative samples. We excluded FE from this RQ, as the total number of available FE samples in our training set is already below 500 (see Section 3.3.6), making it inherently a low-resource task. The experiment results are shown in Table 8.
@@ 4.3.1. Experiment Setup
Model #Training Samples Method CC CM DC
@@ 4.3.2. Results of RQ3
CC You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”.
@@
  • * Prompt 1 shown in this table was curated by the authors. Prompt 2 and Prompt 3, which were generated by ChatGPT, are available in our replication package (replicationPackage, ).
4.4.1. Experiment Setup
@@
In this RQ, we established both heuristic-based (i.e., DesigniteJava and PMD) and best-in-class DL detectors (i.e., DeepSmells (ho2023fusion, ) and AE-Dense (sharma2021code, )) as baselines for comparison. Additionally, In-Context Learning (ICL) (tom2020language, ) has demonstrated impressive performance in various SE tasks, such as code generation (gao2024what, ; jia2024acecoder, ). To evaluate its potential in code smell detection, we explored the effectiveness of two recent and powerful general-purpose LLMs, GPT-4o-mini (hurst2024gpt4o, ) and DeepSeek-v3 (liu2024deepseek, ), under both zero-shot and few-shot settings. We selected these two models because they represent the latest advancements in commercial and open-weight general-purpose LLMs, respectively, and have demonstrated strong performance across a range of natural language and programming-related tasks (rasheed2024task, ; bruni2025benchmarking, ; liu2024deepseek, ; jahin2025unveiling, ).
Specifically, we first designed a prompt template (i.e., prompt 1) for each of the four types of code smells examined in this study, incorporating both instructional text and smell definitions, as shown in Table 9. We then used ChatGPT (ChatGPT, ) to generate two additional variants for each template (i.e., prompt 2 and prompt 3), resulting in three distinct prompts per code smell type. For the zero-shot setting, these prompts were directly provided to GPT-4o-mini and DeepSeek-v3 without any labeled examples. In the few-shot setting, example pairs were randomly sampled from the training set, where each pair consisted of one code sample with a code smell and one without, along with their corresponding labels. These examples were combined with the base prompts to form few-shot configurations containing one, two, or three example pairs. Due to resource constraints (e.g., limited API usage budget), we restricted the few-shot setting to include no more than three example pairs per prompt. As a result, each code smell type yields a total of 12 unique prompt-example configurations, derived from three prompt variants combined with four levels of contextual information (zero-shot, plus one-, two-, and three-pair few-shot settings).
This ICL setup ensures diversity in both prompt formulations and example composition, allowing for a more comprehensive evaluation of general-purpose LLMs and enabling a fair comparison with our proposed PEFT-based methods in the context of code smell detection. Figure 7 presents the performance of GPT-4o-mini and DeepSeek-v3 on code smell detection in both zero- and few-shot settings. Table 10 summarizes the final results for RQ4, where we report only the best-performing configurations for both GPT-4o-mini and DeepSeek-v3 under ICL across all prompts and example variants.
4.4.2. ICL
It can be observed from Figure 7 that DeepSeek-v3 generally demonstrates better effectiveness in code smell detection compared to GPT-4o-mini. Additionally, both models achieve better performance under few-shot settings than in the zero-shot setting. However, their performance does not consistently improve as the number of example pairs increases.
4.4.3. Comparison with State-of-the-Art Approaches
@@
Based on the results presented in Table 10, we can find that LMs fine-tuned with PEFT techniques consistently demonstrate the highest effectiveness among all existing methods for code smell detection. Additionally, among conventional baseline approaches, DeepSeek-v3 achieves the most competitive overall performance. While it shows a significant performance gap compared to LMs tuned with PEFT methods for CM and FE detection (with MCC differences exceeding 10%), its performance in CC and DC detection remains relatively strong, with only a 4%-5% reduction in MCC values, demonstrating its relative strength in these specific code smell types.
Furthermore, our results indicate that traditional DL-based methods (DeepSmells and AE-Dense) exhibit substantially inferior performance compared to both PEFT and the other baseline approaches. This finding underscores the limitations of conventional DL models in code smell detection, particularly when compared to more advanced techniques like PEFT techniques and ICL.
@@ 5. Discussion
Approaches with improved effectiveness for CC code smell detection are needed: According to Fowler (fowler1999refactoring, ), four factors contribute to making programs difficult to work with, one of which is Complex Conditional (CC) logic that complicates modification. The presence of CC in source code hinders testing and maintenance (fowler1999refactoring, ; liu2014automated, ; alhefdhi2022framework, ). Thus, identifying CC code smells in software can significantly improve code quality (liu2014automated, ). However, the effectiveness of existing methods for detecting this type of code smell is considerably limited. As shown in Table 10, existing methods, including heuristic-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells) and AE-Dense), and general-purpose LLMs under ICL settings (e.g., GPT-4o-mini and DeepSeek-v3), all demonstrate limited success in detecting CC, with all MCC values below 50%, and some even as low as 6.63%. Although our proposed method, StarCoderBase-3B with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 74.68% and a MCC value of 49.96%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability. Accurate identification of CC code smells can help developers better refactor the code, ultimately improving the maintainability of the software system. Despite the progress made, the detection of CC remains a challenging task that requires further improvements.
@@
PEFT methods can enhance the capacity of LMs in class-level code understanding: In this study, we selected four representative types of code smells, i.e., CC, CM, FE, and DC, to evaluate the effectiveness of PEFT methods. While only DC is formally categorized as a class-level smell, FE detection also inherently requires class-level analysis as well, as it involves analyzing method-class dependencies (see Section 3.3.1 and Section 3.3.4). According to the recent studies (silva2024detecting, ; mesbah2025leveraging, ), LLMs not specifically fine-tuned for code smell detection generally perform worse when tasked with identifying smells that demand class-level understanding. This aligns with our intuitive expectation that class-level code understanding is more difficult, as it requires capturing higher-level structural relationships (e.g., field-method interactions, class cohesion, and design-level abstractions). Despite this challenge, most existing research on PEFT methods has predominantly concentrated on the their perfomance for method-level code understanding (liu2023empirical, ; wang2023one, ; liu2024delving, ), leaving their effectiveness on class-level tasks largely unexplored. Our work bridges this critical gap by evaluating PEFT methods on code smell detection, a task spans both method- and class-level analysis. Surprisingly, as discussed in Section 4.1.5, our results reveal that LLMs with PEFT methods not only overcome the anticipated class-level challenges but even outperform method-level detection in some cases. These findings suggest that PEFT methods not only improve parameter efficiency but also strengthen the proficiency of LLMs to perform robust code understanding across different levels of granularity.
@@
While larger LMs with PEFT methods often achieve better effectiveness in code smell detection, SLMs still retain distinct advantages: The results of RQ1 reveal that LLMs generally outperform SLMs in code smell detection, particularly in identifying FE. This performance advantage aligns with established neural scaling laws (kaplan2020scaling, ), which posit that model capabilities improve predictably with increased scale. However, SLMs are able to achieve comparable performance in detecting CC and CM, and even significantly outperform LLMs in the detection of DC. Moreover, as illustrated in Figure 5, SLMs exhibit substantially lower peak GPU memory consumption compared to LLMs, making them a more resource-efficient choice. These findings highlight a critical trade-off: while LLMs tend to yield higher absolute performance, SLMs offer a competitive and cost-effective alternative, especially in settings with limited computational resources.
@@
Start model selection for code smell detection with a smaller dataset to optimize resource usage: In low-resource scenarios, even when the dataset size was drastically reduced, the models’ effectiveness did not degrade significantly, and they still yielded promising results, as shown in the results of RQ3. This finding indicates that, even with a significantly smaller dataset, both small and large LMs can maintain robust performance for the code smell detection task. Consequently, beginning fine-tuning experiments with a smaller-size dataset can be an effective strategy to optimize resource usage. This approach allows researchers and practitioners to initially evaluate a wide range of LMs in a shorter time and wither fewer computational resources, saving substantial effort. It enables them to identify the most suitable model for the specific task through an initial evaluation before committing to the more resource-intensive process of training on larger datasets. Additionally, it helps in assessing the trade-offs between resource consumption and performance, providing insight into the models’ scalability. By leveraging smaller datasets for initial evaluation, one can make informed decisions about which models are likely to perform well with limited resources, and further fine-tune or scale up as necessary, without wasting valuable time or computational power.
@@
ICL shows promise as an effective alternative to PEFT methods for certain types of code smells: In RQ4, we investigate the performance of general-purpose LLMs in code smell detection under In-Context Learning (ICL) settings. As shown in Table 10, for CC and DC, the optimal ICL configuration of DeepSeek-v3 achieves MCC scores only 4%–5% lower than those attained by the top-performing PEFT-tuned LMs. This suggests that, for certain types of code smells, well-crafted prompts and a few in-context examples may enable general-purpose LLMs to produce competitive results. Moreover, applying ICL technique requires no model training or parameter updates (zhou2024large, ; tom2020language, ; li2025large, ), which is particularly appealing for resource-constrained scenarios or when model weights cannot be modified (e.g., in commercial LLM APIs). However, for CM and FE, ICL lags significantly behind, with performance deficits of up to 13.69% and 29.25% in MCC, respectively, compared to our PEFT-based approaches. These results indicate that detecting these smell types likely requires deeper model adaptation to capture subtle semantic patterns and structural nuances that few-shot prompting alone cannot convey. Overall, while ICL offers a lightweight and flexible alternative, especially for rapid prototyping or scenarios with limited computational resources, PEFT techniques remain the more robust and reliable solution for achieving high-accuracy code smell detection across a broad range of smell types.
@@
Future work could explore the use of Retrieval-Augmented Generation (RAG) for code smell detection: According to Figure 7, while models generally perform better in code smell detection when provided with randomly selected examples compared to the zero-shot setting, their performance does not consistently improve as the number of example pairs increases. This observation suggests that simply increasing the number of in-context examples without considering their quality, relevance, or diversity may not be sufficient to enhance model effectiveness for the task of code smell detection. By contrast, Retrieval-Augmented Generation (RAG) (lewis2020retrieval, ) can dynamically retrieve the most relevant examples from an external knowledge base to augment the prompts. By grounding predictions in high-quality and contextually appropriate evidence, RAG has the potential to enhance detection accuracy while avoiding the need for fine-tuning or complex prompt engineering. As such, future research could explore whether RAG-based approaches can provide a more effective and scalable solution to PEFT methods for code smell detection.

### v3

@@ 5.1.1. SLMs
Overall, GraphCodeBERT proved the most effective in identifying CC and DC, while UnixCoder excelled at detecting CM and FE. In comparison, CodeT5 fine-tuned with PEFT methods performed the worst across all types of code smells. CodeBERT performed similarly to GraphCodeBERT, albeit slightly worse. For example, in CC detection, whether using full fine-tuning or PEFT methods, CodeBERT consistently achieved lower F1-scores and MCC values compared to GraphCodeBERT, with differences ranging from 0.11% to 5.93% and 0.21% to 12.10%, respectively. This is reasonable because GraphCodeBERT is built upon CodeBERT and further incorporates code structure information (i.e., data flow) during the pre-training phase (guo2021graphcodebert). Across all SLMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods in the code smell detection task, which requires both method- and class-level code understanding. When fine-tuned with PEFT methods, the MCC values of CodeBERT even became negative, reaching -4.05% for LoRA on CC and -2.66% for $(IA)^{3}$ on DC. This suggests that CodeT5 with PEFT methods struggles to perform effectively in code smell detection, highlighting the limitations of its PEFT adaptation. In contrast, the fully fine-tuned CodeT5 demonstrated acceptable performance in code smell detection. The stark performance gap between CodeT5 with full fine-tuning and PEFT methods could stem from CodeT5’s architecture, which may depend more on full fine-tuning for optimal adaptation to the method-level code smell detection task. Compared to full fine-tuning, PEFT methods like LoRA and $(IA)^{3}$ might not provide sufficient adjustments to effectively adapt CodeT5’s encoder-decoder architecture.
@@ 5.1.2. LLMs
Figure 6. Peak GPU memory usage for detecting $C$: (a) SLMs with PEFT methods and Full Fine-Tuning; (b) LLMs with PEFT methods
@@
The detection of CC and CM requires method-level code understanding, whereas identifying FE and DC necessitates class-level analysis (see Section 3.1). By comparing the detection results of CC and CM with those of FE and DC, we observe that LMs demonstrate comparable effectiveness at both levels of code understanding. Notably, the models even achieve better performance on class-level code smells, as evidenced by the fact that the best-performing model yields significantly lower results on CC compared to the other smell types.
@@ 5.1.5. Best PEFT Method
Generally, for SLMs, both $(IA)^{3}$ and prefix tuning achieve strong performance across different models, while for LLMs, $(IA)^{3}$ consistently delivers the best results across all four types of code smells. To be specific, the optimal PEFT method for SLMs on the task of code smell detection depends on the specific model: for instance, CodeBERT performs best with $(IA)^{3}$, whereas UnixCoder tends to favor prefix tuning. In contrast, $(IA)^{3}$ outperforms other PEFT techniques for all LLMs studied, including StarCoderBase Models, Deepseek-Coder Models, and CodeLlama-7B. Additionally, prompt tuning can serve as an alternative for LLMs, though its effectiveness is generally lower. Notably, despite achieving promising results on certain models and code smell types, LoRA exhibits instability in performance and can yield markedly poor effectiveness in some cases. For example, on StarCoderBase-1B for CM detection, LoRA’s MCC drops to 7.01%, while the other three PEFT methods achieve a minimum MCC of 54.01%, resulting in a disparity of over 40%.
As shown in Tables 6 and 7, the additional parameters injected by all PEFT methods are negligible relative to the total model size. Prefix tuning and LoRA generally introduce more trainable parameters than prompt tuning and $(IA)^{3}$, with LoRA adding the most and prompt tuning the least. For SLMs, the absolute differences in the number of added parameters remain modest due to their limited model size, whereas for LLMs the gaps become substantially larger.
Overall, $(IA)^{3}$ introduces a relatively small amount of trainable parameters while delivering strong performance gains for both SLMs and LLMs in code smell detection. For SLMs, prefix tuning still remains a practical alternative; however, for LLMs it brings a considerably larger parameter overhead without providing comparable performance benefits.
@@
According to Tables 6 and 7, all four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for code smell detection. For SLMs (excluding CodeT5), prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ generally achieve varying degrees of improvement in F1-scores and MCC values compared to full fine-tuning across most code smell types. However, for FE, all PEFT methods perform significantly worse than full fine-tuning.
The “#Params” columns in Tables 6 and 7 further show that all PEFT methods introduce a relatively small number of trainable parameters compared to full fine-tuning, typically ranging from tens to hundreds of millions. This demonstrates that, in most cases, PEFT approaches can match or even surpass the performance of standard full fine-tuning for code smell detection while updating far fewer parameters, highlighting their efficiency for the task of code smell detection.
In addition to reducing trainable parameters, PEFT methods also lower computational overhead, particularly peak GPU memory usage, without sacrificing performance. Taking the detection of CC as an example, Figure 6(a) depicts the peak GPU memory consumption during the fine-tuning of the four SLMs. It can be observed that, juxtaposed with full fine-tuning, all PEFT methods lead to a substantial reduction in peak GPU memory usage, which further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
Answer to RQ1: Overall, LLMs perform slightly better than SLMs in code smell detection, except for DC. Among the PEFT methods, both $(IA)^{3}$ and prefix tuning are highly effective for SLMs, whereas $(IA)^{3}$ consistently delivers the best results for LLMs. All PEFT techniques either exceed or match full fine-tuning while requiring far fewer trainable parameters and significantly reducing peak GPU memory usage, underscoring their efficiency for code smell detection.
@@ 5.2. RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
Figure 7. MCC results of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA for code smell detection under different rank settings: (a) $C$; (b) CM; (c) FE; (d) DC
@@ 5.3. RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
To simulate low-resource scenarios, following previous studies (wang2023one; liu2023empirical), we randomly selected 50, 100, 250, and 500 samples for CC, CM, and DC, respectively, based on the size of our dataset. Note that each subset is balanced, containing an equal number of positive and negative samples. We excluded FE from this RQ, as the total number of available FE samples in our training set is already below 500 (see Section 3.6), making it inherently a low-resource task. The experiment results are shown in Table 8.
@@
Model #Training Samples Method $C$ CM DC
@@ 5.3.1. Results of RQ3
$C$ You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”.
@@
  • * Prompt 1 shown in this table was curated by the authors. Prompt 2 and Prompt 3, which were generated by ChatGPT, are available in our replication package (replicationPackage).
@@
Figure 9. Results of DeepSeek-v3 and GPT-4o-mini on detecting CC, CM, FE, and DC with and without labeling criteria
To answer this RQ, we benchmark the performance of our PEFT-based methods against the three categories of baselines established in Section 4.4: heuristics-based detectors (DesigniteJava, PMD), DL-based approaches (DeepSmells, AE-Dense), and general-purpose LLMs (GPT-4o-mini, DeepSeek-v3) with In-Context Learning (ICL).
For the general-purpose LLMs, we evaluated their effectiveness under the ICL paradigm. To ensure the robustness of our evaluation against prompt sensitivity, we designed three prompt formulations for each code smell type: a base prompt containing instructional text and definitions (Prompt 1, as shown in Table 9), and two additional variants (Prompts 2 and 3) generated using ChatGPT (ChatGPT) to introduce linguistic diversity.
These prompts were applied in three ICL settings:
  1. (1)
    Zero-shot setting: prompts used directly without examples;
  2. (2)
    Few-shot setting: prompts augmented with one, two, or three randomly sampled example pairs (consisting of one smelly and one clean code sample) to provide contextual guidance;
  3. (3)
    Criteria-informed setting: prompts enriched with the explicit labeling rules derived from our manual review process (Section 3.5).
This multi-faceted design ensures diversity in both prompt formulation and contextual information, thereby supporting a comprehensive evaluation of general-purpose LLMs and enabling a fair comparison with our PEFT-based methods in the context of code smell detection. In total, this yielded 15 unique configurations per smell type (combinations of three prompt variants and five varying contextual levels).
Figure 8 visualizes the performance trends of GPT-4o-mini and DeepSeek-v3 across the zero- and few-shot settings, while Figure 9 illustrates the impact of incorporating labeling criteria. Table 10 summarizes the final comparative results, reporting the best-performing configuration of each LLM alongside our PEFT methods and other baselines.
5.4.1. Performance of General-Purpose LLMs with ICL
It can be observed from Figures 8 and 9 that DeepSeek-v3 generally demonstrates better effectiveness in code smell detection compared to GPT-4o-mini across all ICL settings. Both models perform better under few-shot settings than in the zero-shot setting; however, their performance does not consistently improve as the number of example pairs increases. Moreover, GPT-4o-mini shows a noticeable performance gain under the criteria-informed setting, whereas DeepSeek-v3 performs better when labeling criteria are not provided. This difference may stem from model characteristics: GPT-4o-mini, a smaller commercial model with stronger instruction alignment, benefits more from explicit and structured guidance such as labeling criteria, which compensates for its limited parameter capacity and contextual reasoning ability. In contrast, DeepSeek-v3, with its larger scale and broader training on code-related corpora, can generalize more effectively from implicit semantics and example patterns, exhibiting stronger reasoning ability even without explicit rules.
5.4.2. Overall Comparison with State-of-the-Art Approaches
@@
Based on the results presented in Table 10, we can find that LMs fine-tuned with PEFT techniques consistently demonstrate the highest effectiveness among all existing methods for code smell detection. Specifically, compared to the best-performing baseline, PEFT-tuned LMs achieve MCC improvements of 5.85%, 13.69%, 12.79%, and 0.33% for CC, CM, FE, and DC, respectively. Among various baseline approaches, DeepSeek-v3 achieves the most competitive overall performance. However, despite its relative strength in detecting CC and DC, a significant performance gap remains: DeepSeek-v3 lags considerably behind PEFT-tuned LMs in CM and FE detection, where the performance differences exceed 10%.
Furthermore, our results indicate that traditional DL-based methods (DeepSmells and AE-Dense) exhibit substantially inferior performance compared to both PEFT and the other baseline approaches. This finding underscores the limitations of conventional DL models in code smell detection, particularly when compared to more advanced techniques like PEFT and ICL.
@@ 6. Discussion
Approaches with improved effectiveness for CC code smell detection are needed: According to Fowler (fowler1999refactoring), four factors contribute to making programs difficult to work with, one of which is Complex Conditional (CC) logic that complicates modification. The presence of CC in source code hinders testing and maintenance (fowler1999refactoring; liu2014automated; alhefdhi2022framework). Thus, identifying CC code smells in software can significantly improve code quality (liu2014automated). However, the effectiveness of existing methods for detecting this type of code smell is considerably limited. As shown in Table 10, existing methods, including heuristics-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells) and AE-Dense), and general-purpose LLMs under ICL settings (e.g., GPT-4o-mini and DeepSeek-v3), all demonstrate limited success in detecting CC, with all MCC values below 50%, and some even as low as 6.63%. Although our proposed method, StarCoderBase-3B with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 74.68% and a MCC value of 49.96%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability. Accurate identification of CC code smells can help developers better refactor the code, ultimately improving the maintainability of the software system. Despite the progress made, the detection of CC remains a challenging task that requires further improvements.
@@
PEFT methods can enhance the capacity of LMs in class-level code understanding: In this study, we selected four representative types of code smells, i.e., CC, CM, FE, and DC, to evaluate the effectiveness of PEFT methods. While only DC is formally categorized as a class-level smell, FE detection also inherently requires class-level analysis as well, as it involves analyzing method-class dependencies (see Section 3.1 and Section 3.4). According to the recent studies (silva2024detecting; mesbah2025leveraging), LLMs not specifically fine-tuned for code smell detection generally perform worse when tasked with identifying smells that demand class-level understanding. This aligns with our intuitive expectation that class-level code understanding is more difficult, as it requires capturing higher-level structural relationships (e.g., field-method interactions, class cohesion, and design-level abstractions). Despite this challenge, most existing research on PEFT methods has predominantly concentrated on the their perfomance for method-level code understanding (liu2023empirical; wang2023one; liu2024delving), leaving their effectiveness on class-level tasks largely unexplored. Our work bridges this critical gap by evaluating PEFT methods on code smell detection, a task spans both method- and class-level analysis. Surprisingly, as discussed in Section 5.1.4, our results reveal that LLMs with PEFT methods not only overcome the anticipated class-level challenges but even outperform method-level detection in some cases. These findings suggest that PEFT methods not only improve parameter efficiency but also strengthen the proficiency of LLMs to perform robust code understanding across different levels of granularity.
When fine-tuning LMs for code smell detection, more trainable parameters do not guarantee better results: Our findings indicate that the effectiveness of fine-tuning is not primarily determined by the number of parameters being updated. Although full fine-tuning adjusts all model parameters, it does not consistently outperform PEFT methods in code smell detection. For many SLMs and most code smell types, PEFT approaches achieve comparable or even superior performance while adding only a small fraction of parameters. This indicates that large-scale parameter updates are not necessary for effective model adaptation in this task. A similar pattern emerges when comparing individual PEFT techniques. LoRA and prefix tuning introduce far more trainable parameters than prompt tuning or $(IA)^{3}$, yet their additional capacity does not reliably translate into better performance. In some cases, LoRA even performs markedly worse despite being the most parameter-intensive PEFT method. In contrast, $(IA)^{3}$ frequently provides the best overall performance while relying on relatively few added parameters. Taken together, these observations confirm that increasing the number of trainable parameters alone does not guarantee better outcomes. Instead, the effectiveness of fine-tuning depends on how efficiently the model’s internal representations are adapted and on the strategic placement of parameter updates for code smell detection.
@@
While larger LMs with PEFT methods often achieve better effectiveness in code smell detection, SLMs still retain distinct advantages: The results of RQ1 reveal that LLMs generally outperform SLMs in code smell detection, particularly in identifying FE. This performance advantage aligns with established neural scaling laws (kaplan2020scaling), which posit that model capabilities improve predictably with increased scale. However, SLMs are able to achieve comparable performance in detecting CC and CM, and even significantly outperform LLMs in the detection of DC. Moreover, as illustrated in Figure 6, SLMs exhibit substantially lower peak GPU memory consumption compared to LLMs, making them a more resource-efficient choice. These findings highlight a critical trade-off: while LLMs tend to yield higher absolute performance, SLMs offer a competitive and cost-effective alternative, especially in settings with limited computational resources.
@@
Start model selection for code smell detection with a smaller dataset to optimize resource usage: In low-resource scenarios, even when the dataset size was drastically reduced, the models’ effectiveness did not degrade significantly, and they still yielded promising results, as shown in the results of RQ3. This finding indicates that, even with a significantly smaller dataset, both small and large LMs can maintain robust performance for the code smell detection task. Consequently, beginning fine-tuning experiments with a smaller-size dataset can be an effective strategy to optimize resource usage. This approach allows researchers and practitioners to initially evaluate a wide range of LMs in a shorter time and with fewer computational resources, saving substantial effort. It enables them to identify the most suitable model for the specific task through an initial evaluation before committing to the more resource-intensive process of training on larger datasets. Additionally, it helps in assessing the trade-offs between resource consumption and performance, providing insight into the models’ scalability. By leveraging smaller datasets for initial evaluation, one can make informed decisions about which models are likely to perform well with limited resources, and further fine-tune or scale up as necessary, without wasting valuable time or computational power.
@@
ICL shows promise as an effective alternative to PEFT methods for certain types of code smells: In RQ4, we investigated the performance of general-purpose LLMs in code smell detection under In-Context Learning (ICL) settings, including zero-shot, few-shot, and criteria-informed setups. As shown in Table 10, for CC and DC, the optimal ICL configuration of DeepSeek-v3 achieves MCC scores only 1%–5% lower than those attained by the top-performing PEFT-tuned LMs. This suggests that, for certain types of code smells, well-crafted prompts with ICL may enable general-purpose LLMs to produce competitive results. Moreover, applying ICL technique requires no model training or parameter updates (zhou2024large; tom2020language; li2025large), which is particularly appealing for resource-constrained scenarios or when model weights cannot be modified (e.g., in commercial LLM APIs). However, for CM and FE, ICL lags significantly behind, with performance deficits of up to 13.69% and 16.62% in MCC, respectively, compared to our PEFT-based approaches. These results indicate that detecting these smell types likely requires deeper model adaptation to capture subtle semantic patterns and structural nuances that few-shot prompting alone cannot convey. Overall, while ICL offers a lightweight and flexible alternative, especially for rapid prototyping or scenarios with limited computational resources, PEFT techniques remain the more robust and reliable solution for achieving high-accuracy code smell detection across a broad range of smell types.
@@
Future work could explore the use of Retrieval-Augmented Generation (RAG) for code smell detection: According to Figure 8, while models generally perform better in code smell detection when provided with randomly selected examples compared to the zero-shot setting, their performance does not consistently improve as the number of example pairs increases. This observation suggests that simply increasing the number of in-context examples without considering their quality, relevance, or diversity may not be sufficient to enhance model effectiveness for the task of code smell detection. By contrast, Retrieval-Augmented Generation (RAG) (lewis2020retrieval) can dynamically retrieve the most relevant examples from an external knowledge base to augment the prompts. By grounding predictions in high-quality and contextually appropriate evidence, RAG has the potential to enhance detection accuracy while avoiding the need for fine-tuning or complex prompt engineering. As such, future research could explore whether RAG-based approaches can provide a more effective and scalable solution to PEFT methods for code smell detection.